OpenStudy (anonymous):

Find the area of a rectangular object which has a length of 5 square root of 8 inches and a width of 2 square root of 5 inches.

OpenStudy (anonymous):

Think of it this way, find both your sqaure roots.

OpenStudy (anonymous):

3.3 is what i get

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

i don't think my teacher wants that format...

OpenStudy (anonymous):

omg i got it!!! Multiply 5√8 by 2√5 to get... 10√(8 * 5) = 10 * 2√(2 * 5) = 20√10 in²

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

Good job!